OpenStudy (anonymous):

First derivative with steps: 1+(2+x)(e^(x/3))

OpenStudy (anonymous):

looks like you need to use the product rule mixed with the quotient rule. Are you familiar with either of these?

OpenStudy (anonymous):

Yes, I already came up with an answer, but I don't think it is right. I need steps done so I can compare them with mine and find the mistake.

OpenStudy (anonymous):

post ur answer

OpenStudy (anonymous):

Ok give me just a second

OpenStudy (anonymous):

1/3(e^(x/3))(x+5)

OpenStudy (anonymous):

\[1+(2+x)(e^{(x/3)})\]\[(2+x)e^{x/3}\frac{1}{3}+x^{x/3}\]

OpenStudy (anonymous):

\[y=uv\]\[y'=uv'+u'v\]

OpenStudy (anonymous):

So is that the answer orate hose steps?

OpenStudy (anonymous):

*or are those

OpenStudy (anonymous):

answer. Not that much steps involved. derivative of (x+2) is 1 derivative of 1 is 0 derivative of e^(x/3) is [e^(x/3)] (1/3)

OpenStudy (anonymous):

so just apply the product rule

OpenStudy (anonymous):

And you are absolutely positive?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

do you think that you would distribute the e^x/3 first before taking the derivative?

OpenStudy (anonymous):

@Hero double check my sol'n plz

OpenStudy (anonymous):

wolfram alpha matches with your answer

OpenStudy (anonymous):

I'm not sure, that's why I asked the question

OpenStudy (anonymous):

I'm going to simplify mine and maybe i'll get urs. hold on plz

OpenStudy (anonymous):

\[(2+x)e^{x/3}\frac{1}{3}+x^{x/3}\]\[\frac{2}{3}e^{x/3}+(1/3)xe^{x/3}+e^{x/3}\]\[e^{x/3}(\frac{5}{3}+(1/3)x)\]\[\frac{1}{3}e^{x/3}(5+x)\]Ya same answer. Sorry if that confused you

OpenStudy (anonymous):

correction on first line: last term should be e^(x/3)